A physics professor demonstrates the Doppler effect by tying a 558.92Hz sound generator to a 1.69-m-long rope and whirling it around her head in a horizontal circle at 97.49 rpm. Assume the room temperature is 10.93oC. What is the highest frequency heard by a student in the classroom?
Answer #1

w = 97.49 rpm = 10.2 rad/s

v = r*w = 1.69 * 10.2 = 17.25 m/s

students will hear the highest frequency when generator approach to the students

f' = fo*(v+vo/v-vs)

here vo = 0 m/s observer are students and they are stationary

v = speed of sound at 10.93 C = 337.864 m/s

f' = 588.998 = 589 Hz